# Hasan Fallahgoul

### v2

@@ 1 Introduction
The integration of machine learning methods into financial prediction has emerged as one of the most active areas of research in empirical asset pricing (Kelly et al. 2024, Gu et al. 2020, Bianchi et al. 2021, Chen et al. 2024, Feng et al. 2020). The appeal is clear: while financial markets generate increasingly high-dimensional data, traditional econometric methods remain constrained by limited sample sizes and the curse of dimensionality. Machine learning promises to uncover predictive relationships that elude traditional linear models by leveraging nonlinear approximations and high-dimensional overparameterized representations, thereby expanding the frontier of return predictability and portfolio construction.
@@
Second, I derive sharp sample complexity bounds that characterize the information-theoretic limits of high-dimensional learning in financial settings. Using PAC-learning theory, [^1] In PAC-learning (Valiant 1984), a predictor is “probably approximately correct” if, with $T\!\gtrsim\!(\text{capacity})/\varepsilon^{2}$ samples, its risk is within $\varepsilon$ of optimal with probability $1-\delta$; I apply these bounds (see Kearns & Vazirani 1994) to gauge when weak return signals are learnable. I establish both exponential and polynomial lower bounds showing when reliable extraction of weak predictive signals becomes impossible regardless of the sophistication of the employed method. These bounds reveal that learning over function spaces with thousands of parameters suggests that reliable learning may require stronger conditions than typically available in typical financial applications. For example, methods claiming to harness 12,000 parameters with 12 monthly observations require signal-to-noise ratios exceeding realistic bounds by orders of magnitude, suggesting that predictive success may arise through mechanisms that differ from the theoretical framework.
@@ 1.1 Literature Review
The remainder of the paper proceeds as follows. Section 2 establishes the theoretical framework and formalizes the theory-practice disconnect in RFF implementations. Section 3 proves that within-sample standardization fundamentally breaks kernel approximation, explaining why claimed theoretical properties cannot hold in practice. Section 4 establishes information-theoretic barriers to high-dimensional learning, showing that genuine complexity benefits are impossible under realistic financial conditions. Section 5 provides numerical validation of the theoretical predictions. Section 6 concludes. All technical details are relegated to a supplementary document containing Appendices A, B, and C.
2 Background and Framework
@@
This section establishes the theoretical framework for analyzing high-dimensional prediction methods in finance. I first formalize the return prediction problem, then examine the critical disconnect between RFF theory and practical implementation that underlies my main results.
@@ 2.1 The Financial Prediction Problem
Consider the fundamental challenge of predicting asset returns using high-dimensional predictor information. I observe predictor vectors $x_{t}\in\mathbb{R}^{K}$ and subsequent returns $r_{t+1}\in\mathbb{R}$ for $t=1,\ldots,T$, with the goal of learning a predictor $\hat{f}:\mathbb{R}^{K}\to\mathbb{R}$ that minimizes expected squared loss $\mathbb{E}[(r_{t+1}-\hat{f}(x_{t}))^{2}]$.
The challenge lies in the fundamental characteristics of financial prediction: signals are weak relative to noise, predictors exhibit complex persistence patterns, and available sample sizes are limited by the nonstationarity of financial markets. These features create a particularly demanding environment for high-dimensional learning methods.
I formalize this environment through three core assumptions that capture the essential features while maintaining sufficient generality for my theoretical analysis.
@@
This assumption captures the essential features of financial prediction that distinguish it from typical machine learning applications. The bounded signal condition and weak SNR scaling reflect the empirical reality that financial predictors typically explain only 1-5% of return variation (Welch & Goyal 2008). The persistence in predictors (eigenvalues of $\Phi$ in $(0,1)$) captures the well-documented dynamics of financial variables like dividend yields and interest rate spreads, which proves crucial for understanding why short training windows lead to mechanical pattern matching rather than genuine learning.
@@
This assumption formalizes the RFF methodology as actually implemented in practice, including the crucial standardization step that has not been analyzed in existing theoretical frameworks. The standardization appears in every practical implementation to improve numerical stability, yet as I prove, it fundamentally alters the mathematical properties of the method.
Assumption 2.3 (Regularity Conditions).
[…]
These technical conditions ensure that concentration inequalities apply and that my convergence results hold with high probability. The conditions are mild and satisfied in typical financial applications. [^3] For example, for the Kelly et al. (2024) setup with $K=15$ predictors and $T=12$ training windows, these conditions hold almost surely since continuous economic variables generically satisfy the required independence properties.
Assumption 2.4 (Affine Independence of the Sample).
[…]
This assumption enters my analysis through the small-ball probability estimates needed to establish convergence of standardized kernels. The full-rank requirement ensures that the linear change of variables $(\omega,b)\mapsto(2\omega^{\top}x_{t}+2b)_{t=1}^{T}$ is bi-Lipschitz on bounded sets, enabling geometric control that yields exponential small-ball bounds and finiteness of key expectations. In Kelly et al.’s empirical design with $K=15$ predictors and $T=12$ months, the matrix $A$ is $16\times 12$, and since elements are continuous macroeconomic variables, affine dependence has Lebesgue measure zero, making this assumption mild.
Assumption 2.5 (Sub-Gaussian RFFs).
For every unit vector $u\in\mathbb{R}^{P}$, the scalar $u^{\top}z(x)$ is $\kappa$-sub-Gaussian under $x\sim\mu$: $\;\mathbb{E}[\exp(t\,u^{\top}z(x))]\leq\exp(\tfrac{1}{2}\kappa^{2}t^{2})$ for all $t\in\mathbb{R}$.
Assumption 2.5 requires that linear combinations $u^{\top}z(x)$ of the random Fourier features are sub-Gaussian with parameter $\kappa$, ensuring $\mathbb{E}[\exp(t\,u^{\top}z(x))]\leq\exp(\frac{1}{2}\kappa^{2}t^{2})$ for all unit vectors $u\in\mathbb{R}^{P}$ and scalars $t$. This concentration condition is essential for applying uniform convergence results and obtaining non-asymptotic bounds on the empirical feature covariance matrix that appear in our sample complexity analysis. The assumption is standard in high-dimensional learning theory and is automatically satisfied for RFF with bounded support: since $z_{i}(x)=\sqrt{2}\cos(\omega_{i}^{\top}x+b_{i})\in[-\sqrt{2},\sqrt{2}]$, each feature is bounded, and linear combinations of bounded random variables are sub-Gaussian with parameter $\kappa=O(\sqrt{P})$. This ensures that concentration inequalities apply to the feature covariance estimation, enabling our PAC-learning bounds while remaining satisfied in all practical RFF implementations.
@@ 2.2 The Theory-Practice Disconnect in Random Fourier Features
The foundation of high-dimensional prediction methods in finance rests on RFF theory, yet a fundamental disconnect exists between theoretical guarantees and practical implementation. Understanding this disconnect is crucial for interpreting what these methods actually accomplish.
@@ 2.2.2 What Actually Happens in Practice
$\tilde{z}_{i}(x)=\frac{z_{i}(x)}{\hat{\sigma}_{i}},\quad\hat{\sigma}_{i}^{2}=\frac{1}{T}\sum_{t=1}^{T}z_{i}(x_{t})^{2}$ (2.2)
This standardization fundamentally alters the mathematical properties of the method. The standardized empirical kernel becomes:
@@
This standardized kernel no longer converges to the Gaussian kernel. Instead, as I prove in Theorem 3.1, it converges to a training-set dependent limit $k^{*}_{\text{std}}(x,x^{\prime}|\mathcal{T})\neq k_{G}(x,x^{\prime})$ that violates the shift-invariance and stationarity properties required for kernel methods.

### v3

@@ 1 Introduction
The integration of machine learning methods into financial prediction has emerged as one of the most active areas of research in empirical asset pricing (Kelly et al. 2024, Gu et al. 2020, Bianchi et al. 2021, Chen et al. 2024, Fallahgoul et al. 2024, Feng et al. 2020). The appeal is clear: while financial markets generate increasingly high-dimensional data, traditional econometric methods remain constrained by limited sample sizes and the curse of dimensionality. Machine learning promises to uncover predictive relationships that elude traditional linear models by leveraging nonlinear approximations and high-dimensional overparameterized representations, thereby expanding the frontier of return predictability and portfolio construction.
@@
However, fundamental questions remain about the information-theoretic feasibility of high-dimensional learning under the weak signal conditions typical in finance. While Kelly et al. (2024) demonstrate that complexity can be beneficial under appropriate theoretical conditions, their analysis does not address whether realistic financial environments provide sufficient signal strength and sample sizes to support reliable learning in high-dimensional regimes. This gap is particularly pronounced given the notoriously challenging nature of return prediction, where signals are weak, data are limited, and the signal-to-noise ratios may be insufficient to support the extraction of predictive information from thousands of features using dozens of training observations.
@@
Second, I derive sharp sample complexity bounds that characterize the information-theoretic limits of high-dimensional learning in financial settings. Using PAC-learning theory,111 In PAC-learning (Valiant 1984), a predictor is “probably approximately correct” if, with $T\!\gtrsim\!(\text{capacity})/\varepsilon^{2}$ samples, its risk is within $\varepsilon$ of optimal with probability $1-\delta$; I apply these bounds (see Kearns & Vazirani 1994) to gauge when weak return signals are learnable. I establish both exponential and polynomial lower bounds showing when reliable extraction of weak predictive signals becomes impossible regardless of the sophistication of the employed method. These bounds reveal that learning with thousands of parameters using typical financial sample sizes requires signal-to-noise ratios exceeding realistic bounds by orders of magnitude, demonstrating that the ”virtue of complexity” regime ($P\gg T$) is information-theoretically infeasible under standard financial conditions. This explains why apparent success in high-dimensional regimes must arise through alternative mechanisms rather than genuine complex learning (Nagel 2025).
This capability to flexibly trade computational cost for representational power makes RFF a powerful tool. However, this paper reveals two fundamental barriers to realizing these benefits in practice. First, I prove that the standardization procedures universally employed violate the theoretical conditions required for kernel approximation (Section 4). Second, I establish information-theoretic limits showing that reliable learning in high-dimensional regimes requires sample sizes and signal strengths far exceeding those available in typical financial applications (Section 5).
@@ 1.1 Literature Review
The remainder of the paper proceeds as follows. Section 2 provides background on kernel methods and Random Fourier Features, explaining the theoretical foundations and practical implementations that motivate our analysis. Section 3 establishes the theoretical framework and formalizes the theory-practice disconnect in RFF implementations. Section 4 proves that within-sample standardization fundamentally breaks kernel approximation, explaining why claimed theoretical properties cannot hold in practice. Section 5 establishes information-theoretic barriers to high-dimensional learning, showing that genuine complexity benefits are impossible under realistic financial conditions. Section 6 provides numerical validation of the theoretical predictions. Section 7 concludes. All technical details are relegated to a supplementary document containing Appendices A, B, and C.
2 Kernel Methods and Random Fourier Features
Kernel regression provides a powerful non-parametric framework for capturing complex relationships by implicitly mapping inputs into high-dimensional feature spaces (Schölkopf & Smola 2002). For a given kernel function $k(\cdot,\cdot)$, the solution is constrained by the Representer Theorem to lie within the span of the training data. Predictions thus take the form $\hat{f}(x)=\sum_{t=1}^{T}\alpha_{t}k(x,x_{t})$, where the coefficients $\bm{\alpha}$ are learned. The Gaussian kernel, $k_{G}(x,x^{\prime})=\exp(-\gamma^{2}\|x-x^{\prime}\|^{2}/2)$, is particularly attractive due to its universality and smooth, localized similarity structure.
To prevent overfitting and ensure numerical stability, kernel methods are often paired with ridge regularization, which involves solving the optimization problem $\min_{\bm{\alpha}}\|\mathbf{y}-\mathbf{K}\bm{\alpha}\|^{2}+\lambda\|\bm{\alpha}\|^{2}$. This yields the stable closed-form solution $\bm{\alpha}^{*}=(\mathbf{K}+\lambda\mathbf{I})^{-1}\mathbf{y}$, where $\mathbf{K}$ is the $T\times T$ kernel matrix with entries $\mathbf{K}_{ij}=k(x_{i},x_{j})$. However, kernel methods face a fundamental computational bottleneck: constructing and inverting the kernel matrix requires $O(T^{2})$ storage and $O(T^{3})$ operations, respectively, making them impractical for large datasets.
RFF circumvent this limitation by constructing an explicit, finite-dimensional feature map $z(x)\in\mathbb{R}^{P}$ that approximates the kernel function (Rahimi & Recht 2007). The theoretical foundation relies on Bochner’s theorem, which states that a continuous and shift-invariant kernel $k(x,x^{\prime})=k(x-x^{\prime})$ is the Fourier transform of a non-negative measure $p(\omega)$. This allows the kernel to be expressed as an expectation:
[…]
For the Gaussian kernel, this distribution $p(\omega)$ is a Gaussian $N(0,\gamma^{2}I)$. RFF approximates this expectation using a Monte Carlo estimate with $P$ samples, forming features $z_{i}(x)=\sqrt{2}\cos(\omega_{i}^{T}x+b_{i})$ with $\omega_{i}\sim N(0,\gamma^{2}I)$ and $b_{i}\sim\text{Uniform}[0,2\pi]$, such that $k_{G}(x,x^{\prime})\approx z(x)^{T}z(x^{\prime})$.
The crucial insight is that kernel ridge regression becomes a linear ridge regression problem in the RFF feature space: $\hat{f}(x)=z(x)^{T}\mathbf{w}^{*}$. This approach offers a dual advantage depending on the chosen dimensionality $P$:
  • Computational Speedup ($P\ll T$): For large-scale problems, one can choose a low number of features $P\ll T$ to drastically reduce complexity. The optimal weights $\mathbf{w}^{*}$ are found using the primal solver $\mathbf{w}^{*}=(\mathbf{Z}^{T}\mathbf{Z}+\lambda\mathbf{I})^{-1}\mathbf{Z}^{T}\mathbf{y}$, where $\mathbf{Z}$ is the $T\times P$ feature matrix. The complexity is $O(P^{2}T+P^{3})$, a significant improvement over the original $O(T^{3})$.
  • High-Dimensional Expansion ($P\gg T$): RFF can overcome the representational limit of T basis functions imposed by the Representer Theorem (Schölkopf et al. 2001). By choosing $P\gg T$, we can construct a richer model, a regime termed the ”virtue of complexity” (Kelly et al. 2024). While the primal solver becomes inefficient in this setting, the solution can be found tractably using the equivalent dual form:
    $w^{*}=Z^{\top}(ZZ^{\top}+\lambda I)^{-1}y.$ (2.2)
    This requires inverting a smaller $T\times T$ matrix, making the computation (with complexity $O(T^{2}P)$) feasible while still harnessing the power of a very high-dimensional feature space. However, whether such high-dimensional representations genuinely improve prediction performance under the weak signal conditions of finance is a central question that this paper investigates.
This capability to flexibly trade computational cost for representational power makes RFF a powerful tool. However, as I demonstrate below, the standardization procedures universally employed in practical RFF implementations fundamentally violate the theoretical conditions required for this kernel approximation, creating a critical gap between theory and practice.
2.1 An Illustrative Example
Consider predicting next-month returns using two macro predictors. We have training data:
[…]
$(x_{2},y_{2})$ $=([0.025,0.035],0.020)$ (2.4)
$(x_{3},y_{3})$ $=([0.018,0.028],0.010).$ (2.5)
  • Traditional Kernel Ridge: The $3\times 3$ kernel matrix requires computing $k_{G}(x_{i},x_{j})$ for all pairs, then solving $\bm{\alpha}^{*}=(\mathbf{K}+\lambda\mathbf{I})^{-1}\mathbf{y}$. The method is limited to 3 basis functions.
  • RFF with P = 1000: Generate 1000 random features $z_{i}(x)=\sqrt{2}\cos(\omega_{i}^{T}x+b_{i})$ where $\omega_{i}\sim N(0,\gamma^{2}I)$. This creates a $3\times 1000$ feature matrix $\mathbf{Z}$. Using the dual form $\mathbf{w}^{*}=\mathbf{Z}^{T}(\mathbf{Z}\mathbf{Z}^{T}+\lambda\mathbf{I})^{-1}\mathbf{y}$, we invert only a $3\times 3$ matrix while harnessing 1000-dimensional representations.
  • The Standardization Issue: In practice, features are standardized using the observed training data: $\tilde{z}_{i}(x)=z_{i}(x)/\hat{\sigma}_{i}$ where
    $\hat{\sigma}_{i}^{2}=\frac{1}{3}\sum_{t=1}^{3}z_{i}(x_{t})^{2}=\frac{1}{3}[z_{i}(x_{1})^{2}+z_{i}(x_{2})^{2}+z_{i}(x_{3})^{2}].$ (2.6)
    Since each $\hat{\sigma}_{i}$ depends on the specific observed values $\{x_{1},x_{2},x_{3}\}$ in our training sample, different training windows will produce different standardization factors, even for the same underlying data generating process. [^2] Note that I estimate the feature’s scale using the sample second moment, a valid approach that assumes the feature mean is zero; otherwise, the sample mean would first need to be properly calculated and removed. This training-set dependence fundamentally violates the theoretical requirement that $z_{i}(x)z_{i}(x^{\prime})\to k_{G}(x,x^{\prime})$ regardless of the training data used. The standardized kernel becomes $k^{*}_{std}(x,x^{\prime}|\{x_{1},x_{2},x_{3}\})$ rather than the shift-invariant Gaussian kernel $k_{G}(x,x^{\prime})$, destroying the mathematical foundation that justifies the RFF approximation.
While the high-dimensional expansion regime ($P\gg T$) promises substantial benefits, fundamental questions remain about when such complexity can genuinely improve prediction performance. In financial applications, where signals are weak and training samples are limited, information-theoretic constraints may render the promised gains illusory. This paper provides the first rigorous analysis of these limitations, showing that the combination of weak signal-to-noise ratios typical in finance and moderate sample sizes creates information-theoretic barriers that prevent reliable learning in high-dimensional regimes. Specifically, our minimax analysis reveals that for realistic financial parameters, reliable extraction of predictive signals requires sample sizes far exceeding those typically available, suggesting that apparent success in $P\gg T$ regimes must arise through mechanisms fundamentally different from the theoretical framework.
3 Theoretical Setup and Core Assumptions
@@
This section establishes the theoretical framework for analyzing high-dimensional prediction methods in finance. I first formalize the return prediction problem and core assumptions, then examine the critical disconnect between RFF theory and practical implementation that underlies my main results.
@@ 3.1 The Financial Prediction Problem
Consider the fundamental challenge of predicting asset returns using high-dimensional predictor information. We observe predictor vectors $x_{t}\in\mathbb{R}^{K}$ and subsequent returns $r_{t+1}\in\mathbb{R}$ for $t=1,\ldots,T$, with the goal of learning a predictor $\hat{f}:\mathbb{R}^{K}\to\mathbb{R}$ that minimizes expected squared loss $\mathbb{E}[(r_{t+1}-\hat{f}(x_{t}))^{2}]$.
The challenge lies in the fundamental characteristics of financial prediction: signals are weak relative to noise, predictors exhibit complex persistence patterns, and available sample sizes are limited by the nonstationarity of financial markets. These features create a particularly demanding environment for high-dimensional learning methods, which I formalize through the following core assumptions.
@@
This assumption captures the essential challenge of financial prediction: signals are weak (low $R^{2}$), predictors are persistent (like dividend yields and interest rate spreads), and noise dominates returns. The weak signal-to-noise ratio reflects the empirical reality that financial models typically explain only 1-5% of return variation (Welch & Goyal 2008).
@@
This assumption formalizes the RFF methodology as implemented in practice. It includes the crucial standardization step, which is commonly applied to ensure features have comparable scales for regularization and to improve numerical stability. Despite its ubiquity in practical applications, this step has not been analyzed in existing theoretical frameworks.
My framework, including Assumption 3.2, uses the canonical RFF construction for its clarity and prominence in the literature. It is crucial, however, to analyze the specific variant used in Kelly et al. (2024). Their method generates a pair of features, $[\sin(\gamma\omega_{i}^{\top}x),\cos(\gamma\omega_{i}^{\top}x)]$, for each random projection $\omega_{i}$. I now provide a brief technical motivation for this choice and explain why my main results on the breakdown of kernel approximation remain fully relevant.
The theoretical goal of RFF is to approximate a shift-invariant kernel $k(x-x^{\prime})$, which can be expressed via Bochner’s theorem as the expectation $\mathbb{E}_{\omega}[\cos(\omega^{\top}(x-x^{\prime}))]$. The standard RFF formulation, $z_{i}(x)=\sqrt{2}\cos(\omega_{i}^{\top}x+b_{i})$, achieves this in expectation over the random phase $b_{i}$, as $\mathbb{E}_{b_{i}}[z_{i}(x)z_{i}(x^{\prime})]=\cos(\omega_{i}^{\top}(x-x^{\prime}))$. The variant used by Kelly et al. (2024) achieves this same objective more directly. By the trigonometric identity $\cos(A-B)=\cos(A)\cos(B)+\sin(A)\sin(B)$, the inner product of their feature pair $[\sin(\omega_{i}^{\top}x),\cos(\omega_{i}^{\top}x)]$ with a corresponding pair at $x^{\prime}$ is deterministically equal to $\cos(\omega_{i}^{\top}(x-x^{\prime}))$. This removes the Monte Carlo randomness associated with the phase shift $b_{i}$, potentially leading to a lower-variance approximation of the kernel for a finite number of features $P$, as documented by Sutherland & Schneider (2015).
Despite the elegance of this formulation, its theoretical properties are invalidated by the same practical implementation step we analyze: data-dependent standardization. Kelly et al. (2024) confirm that they standardize each feature by its standard deviation within the training sample. The deterministic relationship described above is therefore broken. The inner product of the standardized features becomes:
$\frac{\sin(\omega_{i}^{\top}x)\sin(\omega_{i}^{\top}x^{\prime})}{\hat{\sigma}_{s,i}^{2}}+\frac{\cos(\omega_{i}^{\top}x)\cos(\omega_{i}^{\top}x^{\prime})}{\hat{\sigma}_{c,i}^{2}}$ (3.1)
where the scaling factors $\hat{\sigma}_{s,i}^{2}$ and $\hat{\sigma}_{c,i}^{2}$ are calculated from the training data. The introduction of these data-dependent denominators fundamentally alters the function being approximated. The resulting kernel is no longer the shift-invariant target but a complex, data-dependent one. Thus, the kernel breakdown theorem, which hinges on this exact mechanism, remains directly applicable and essential for understanding the behavior of the method as implemented.
@@ 3.2 The Theory-Practice Disconnect in Random Fourier Features
Having established the financial prediction environment, we now examine the fundamental disconnect between RFF theory and practical implementation. This disconnect is not merely technical—it undermines the entire theoretical justification for these methods.
@@ 3.2.2 What Actually Happens in Practice
$\tilde{z}_{i}(x)=\frac{z_{i}(x)}{\hat{\sigma}_{i}},\quad\hat{\sigma}_{i}^{2}=\frac{1}{T}\sum_{t=1}^{T}z_{i}(x_{t})^{2}.$ (3.3)
Since each $\hat{\sigma}_{i}$ depends on the specific observed training sample, different training windows produce different standardization factors, even for the same underlying data generating process. This training-set dependence fundamentally violates the theoretical requirement that $z_{i}(x)z_{i}(x^{\prime})\to k_{G}(x,x^{\prime})$ regardless of the training data used. The standardized empirical kernel:
@@
no longer converges to the Gaussian kernel, but instead converges to a training-set dependent limit that violates the mathematical foundations required for kernel methods.
The standardization procedure $\hat{\sigma}_{i}^{2}=\frac{1}{T}\sum_{t=1}^{T}z_{i}(x_{t})^{2}$ creates features whose kernel approximation properties depend on the absolute positions of training points $\{x_{t}\}_{t=1}^{T}$, not just the relative distances $\|x-x^{\prime}\|$ required for shift-invariant kernels. This fundamentally violates the mathematical foundation of RFF theory.
3.3 Technical Conditions
To establish my main theoretical results, I require three additional technical assumptions that ensure proper convergence and concentration properties.
Assumption 3.3 (Regularity Conditions).
The input distribution has bounded support and finite moments, ensuring well-defined feature covariance $\Sigma_{z}=\mathbb{E}[z(x)z(x)^{\top}]$ satisfying $c_{z}I_{P}\preceq\Sigma_{z}\preceq C_{z}I_{P}$ for constants $0<c_{z}\leq C_{z}$. Training samples satisfy standard non-degeneracy conditions.
This is a standard technical requirement in high-dimensional learning theory. The assumption of bounded support is mild and realistic for financial predictors, which naturally operate within a finite range (e.g., interest rates, dividend yields). This well-behaved nature of the inputs ensures the resulting random features also have finite moments. The condition that the feature covariance matrix $\Sigma_{z}$ has eigenvalues bounded away from zero and infinity ($c_{z}I_{P}\preceq\Sigma_{z}\preceq C_{z}I_{P}$) is crucial for mathematical stability. It guarantees the features are not perfectly redundant and have bounded variance, a necessary prerequisite for the concentration inequalities that form the basis of my non-asymptotic proofs.
Assumption 3.4 (Affine Independence of the Sample).
Let $x_{1},\ldots,x_{T}\in\mathbb{R}^{K}$ be the training sample of predictors. The augmented vectors $(x_{t},1)$ are affinely independent. Equivalently, the $(K+1)\times T$ matrix formed by these augmented vectors has full column rank $T$.
This geometric condition on the training data is essential for the analysis of the standardized kernel. It enters our proofs through the small-ball probability estimates required to control the behavior of the random denominators $\hat{\sigma}_{i}^{2}$. The full-rank requirement ensures that the linear map from the random parameters to the feature arguments, $(\omega,b)\mapsto(2\omega^{\top}x_{t}+2b)_{t=1}^{T}$, is locally bi-Lipschitz. This enables the use of powerful geometric tools to establish exponential small-ball bounds, which are critical for proving the convergence of the standardized kernel and the finiteness of key expectations. In typical financial applications where predictors are continuous macroeconomic variables, any exact affine dependence among the training points has Lebesgue measure zero, making this a mild assumption.
Assumption 3.5 (Sub-Gaussian Random Features).
For every unit vector $u\in\mathbb{R}^{P}$, the scalar random variable $u^{\top}z(x)$ is $\kappa$-sub-Gaussian under the data-generating distribution $x\sim\mu$. That is, for all $t\in\mathbb{R}$:
$\mathbb{E}_{\mu,\omega,b}[\exp(t\cdot u^{\top}z(x))]\leq\exp\left(\frac{1}{2}\kappa^{2}t^{2}\right).$ (3.5)
This is a standard concentration assumption in high-dimensional learning theory and is automatically satisfied in all practical RFF implementations. Since each feature $z_{i}(x)=\sqrt{2}\cos(\omega_{i}^{\top}x+b_{i})$ is bounded in $[-\sqrt{2},\sqrt{2}]$, any linear combination $u^{\top}z(x)$ is also bounded. By Hoeffding’s lemma (Hoeffding 1963), any bounded random variable is sub-Gaussian. This condition is essential for applying uniform convergence results (e.g., matrix Bernstein inequalities) to control the difference between the empirical feature covariance matrix and its population counterpart. These non-asymptotic bounds are fundamental to my sample complexity analysis and allow me to derive results that hold with high probability for finite samples.
Having established the fundamental theory-practice disconnect in RFF implementations, the following sections provide rigorous mathematical analysis of its implications. Section 4 proves that standardization fundamentally breaks kernel approximation properties (Theorem 4.1), while Section 5 establishes information-theoretic limits showing that genuine high-dimensional learning is impossible under realistic financial conditions (Theorems 5.1 and 5.2). Together, these results explain why apparent ”virtue of complexity” successes must arise through mechanisms fundamentally different from the theoretical framework.